For a discrete random variable that is poission distributed with a mean i=20 which one of the following statement is incorrect
1. P (X=0)
2. P (X-0 )
3. P (X-1)
4. P (X-7)
5. The variance a is 4
1
Expert's answer
2020-11-03T16:50:01-0500

"X\sim Po(\lambda)"

"P(X=x)=\dfrac{e^{-\lambda}\cdot\lambda^x}{x!}"

"mean=\lambda=Var(x)"

Given "mean=20"

Then "\lambda=20"

1.


"P(X=0)=\dfrac{e^{-20}\cdot20^0}{0!}="

"=e^{-20}\approx0.0000000021\approx0"

2.


"P(X>0)=1-P(X=0)=1-\dfrac{e^{-20}\cdot20^0}{0!}="

"=1-e^{-20}\approx0.9999999979\approx1"

3.


"P(X>1)=1-P(X=0)-P(X=1)="

"=1-\dfrac{e^{-20}\cdot20^0}{0!}-\dfrac{e^{-20}\cdot20^1}{1!}="

"=1-21\cdot e^{-20}\approx0.9999999567\approx1"

4.


"P(X>7)=1-P(X=0)-P(X=1)-"

"-P(X=2)-P(X=3)-P(X=4)-"

"-P(X=5)-P(X=6)-P(X=7)="

"=1-\dfrac{e^{-20}\cdot20^0}{0!}-\dfrac{e^{-20}\cdot20^1}{1!}-\dfrac{e^{-20}\cdot20^2}{2!}-"

"-\dfrac{e^{-20}\cdot20^3}{3!}-\dfrac{e^{-20}\cdot20^4}{4!}-\dfrac{e^{-20}\cdot20^5}{5!}-"

"-\dfrac{e^{-20}\cdot20^6}{6!}-\dfrac{e^{-20}\cdot20^7}{7!}\approx"

"\approx0.9992214099"

5.

"Var(X)=\lambda=20\not=4"